# Random time-shift approximation enables hierarchical Bayesian inference of mechanistic within-host viral dynamics models on large datasets

### v1

@@ 2.7 Inference
The form of Eq. (9) suggests a Gibbs sampler that draws from the conditional distributions for each of the individuals ($i=1,\dots,N$) and the shared parameters $(\boldsymbol{\phi},\kappa)$. However, sampling directly from the conditionals is not possible as the distributions cannot be normalised—or recognised as standard distributions—so we use a Metropolis-Hastings step within the Gibbs sampler. The conditional distributions for the individual parameters, noting that $\boldsymbol{\theta}_{i}$ are independent of $\boldsymbol{\Theta}_{-i}=\left\{\boldsymbol{\theta}_{1},\dots,\boldsymbol{\theta}_{i-1},\boldsymbol{\theta}_{i+1},\dots,\boldsymbol{\theta}_{N}\right\}$ given $\boldsymbol{\phi}$, are given by
$f(\boldsymbol{\theta}_{i}\,|\,\boldsymbol{\phi},\kappa,\mathcal{D})$ $\propto f(\boldsymbol{y}_{i}\,|\,\boldsymbol{\theta}_{i},\kappa)f(\boldsymbol{\theta}_{i}\,|\,\boldsymbol{\phi}).$
The conditional distribution for $\kappa$ and $\phi$ is
$f(\boldsymbol{\phi},\kappa\,|\,\boldsymbol{\Theta},\mathcal{D})\propto f(\kappa)f(\boldsymbol{\phi})\prod_{i=1}^{N}f(\boldsymbol{y}_{i}\,|\,\boldsymbol{\theta}_{i},\kappa)f(\boldsymbol{\theta}_{i}\,|\,\boldsymbol{\phi}).$ (15)
Let the current draw be $\{\boldsymbol{\Theta}^{(k)},\boldsymbol{\phi}^{(k)},\kappa^{(k)}\}$ and we use the notation $\boldsymbol{\Theta}_{a:b}=\{\boldsymbol{\theta}_{a},\dots,\boldsymbol{\theta}_{b}\}$ for integers $a$ and $b$ ($b>a$). Define the following collections
$\psi_{i}^{(k+1)}$ =$=\left\{\boldsymbol{\Theta}_{1:i}^{(k+1)},\boldsymbol{\Theta}_{i+1:N}^{(k)}\right\},\quad i=1,\dots,N-1,$
$\psi_{N}^{(k+1)}$ =$=\left\{\boldsymbol{\Theta}_{1:N}^{(k+1)}\right\}.$
We choose a relatively simple multivariate Gaussian proposal distributions centered on the current values of the parameters being drawn, with $N+1$ separate covariance matrices per step of the method, i.e.
$\boldsymbol{\theta}_{i}^{\prime}|\boldsymbol{\theta}_{i}^{(k)}$ $\sim\mathcal{N}(\boldsymbol{\theta}_{i}^{(k)},\boldsymbol{\Sigma}_{i}),\quad i=1,\dots,N,$
$(\boldsymbol{\phi}^{\prime},\kappa^{\prime})|(\boldsymbol{\phi}^{(k)},\kappa^{(k)})$ $\sim\mathcal{N}((\boldsymbol{\phi}^{(k)},\kappa^{(k)}),\boldsymbol{\Sigma}_{N+1}),$
Blocking the proposals in this way enables the subject specific proposals as well as the proposal for the shared parameters, $\kappa$ and $\boldsymbol{\phi}$, to be tuned separately. The covariance matrices are tuned by performing a sufficiently large pilot run. Denote the PDF of the proposal distributions by $Q(\cdot\,|\,\cdot)$, where the arguments specify the particular proposal distribution being used. The acceptance probabilities for individual parameters ($i=1,\dots,N$) are given by
$\alpha_{1,i}$ $=\min\left\{\frac{f(\boldsymbol{\theta}_{i}^{\prime}\,|\,\psi_{i-1}^{(k+1)},\boldsymbol{\phi}^{(k)},\kappa^{(k)},\mathcal{D})}{f(\boldsymbol{\theta}_{i}^{(k)}\,|\,\psi_{i-1}^{(k+1)},\boldsymbol{\phi}^{(k)},\kappa^{(k)},\mathcal{D})}\times\frac{Q(\boldsymbol{\theta}_{i}^{(k)}\,|\,\boldsymbol{\theta}_{i}^{\prime})}{Q(\boldsymbol{\theta}_{i}^{\prime}\,|\,\boldsymbol{\theta}_{i}^{(k)})},1\right\},$ (16)
Similarly, the acceptance probability for the shared parameters is
$\alpha_{2}$ $=\min\left\{\frac{f(\boldsymbol{\phi}^{\prime},\kappa^{\prime}\,|\,\psi_{N}^{(k+1)},\mathcal{D})}{f(\boldsymbol{\phi}^{(k)},\kappa^{(k)}\,|\,\psi_{N}^{(k+1)},\mathcal{D})}\times\frac{Q(\boldsymbol{\phi}^{(k)},\kappa^{(k)}\,|\,\boldsymbol{\phi}^{\prime},\kappa^{\prime})}{Q(\boldsymbol{\phi}^{\prime},\kappa^{\prime}\,|\,\boldsymbol{\phi}^{(k)},\kappa^{(k)})},1\right\}.$ (17)
Note that the proposal distributions are symmetric and so the ratio of the proposal distributions in Eq. (16) and Eq. (17) are equal to $1$. Algorithm 2 outlines a single step of the Metropolis-within-Gibbs sampler.
Algorithm 2 Single step of the Metropolis-within-Gibbs algorithm
Inputs:
  • Current draw $(\boldsymbol{\Theta}^{(k)},\boldsymbol{\phi}^{(k)},\kappa^{(k)})$;
  • Conditional distributions; $f(\boldsymbol{\theta}_{i}\,|\,\boldsymbol{\phi},\kappa,\mathcal{D})$, $f(\boldsymbol{\phi},\kappa\,|\,\boldsymbol{\Theta},\mathcal{D})$;
  • Transition kernels; $Q(\boldsymbol{\theta}_{i}^{\prime}\,|\,\boldsymbol{\theta}_{i})$ for $i=1,\dots,N$, $Q(\boldsymbol{\phi}^{\prime},\kappa^{\prime}\,|\,\boldsymbol{\phi},\kappa)$
1:for $i=1,\dots,N$ do
2:     Propose individual parameters $\boldsymbol{\theta}_{i}^{\prime}\sim Q(\boldsymbol{\theta}_{i}^{\prime}\,|\,\boldsymbol{\theta}_{i}^{(k)})$.
3:     Compute $\alpha_{1,i}$ (Eq. (16)) and accept proposal with probability $\alpha_{1,i}$.
4:end for
5:Propose shared parameters $\boldsymbol{\phi}^{\prime},\kappa^{\prime}\sim Q(\boldsymbol{\phi}^{\prime},\kappa^{\prime}\,|\,\boldsymbol{\phi}^{(k)},\kappa^{(k)})$.
6:Compute $\alpha_{2}$ (Eq. (17)) and accept proposal with probability $\alpha_{2}$.
2.8 Implementation details
2.8.1 Parallelisation of the inference method
The structure of the inference algorithm and the form of the conditional distributions (Eq. (15)) allow for straightforward parallelisation. The for loop (lines 1-4 of Algorithm 2) samples the VL model parameters $\boldsymbol{\theta}_{i}$ independently of the other individuals given the shared parameters. This means that the sampling of the individual parameters can be parallelised across the individuals. Additionally, parallelisation can also be used when computing the acceptance probability in line 6 as the likelihood contributions from each individual in Eq. (15) (the $\mathcal{L}_{i}(\boldsymbol{\theta}_{i},\kappa)$ terms inside the product) can be computed in parallel conditional on a draw of the hyper-parameters $\boldsymbol{\phi}$.
2.8.2 Amortised learning of parameters for the time-shift distribution
One particular challenge with using the time-shift approach in a statistical inference framework is that computing the parameters of the distribution is itself a computationally expensive task. Each call of the likelihood requires optimisation to be carried out to determine the parameters of the time-shift distribution [17]. To overcome this performance bottleneck, we utilise a neural network (NN) to learn the parameters of the time-shift distribution, $(a,b,p)$, given VL model parameters, $\boldsymbol{\theta}=(R_{0},k,\delta,\rho,c)$. Note the removal of $t_{0}$ from the VL model parameters as the time-shift distribution does not depend on the infection time. This process is referred to as amortised optimisation [48, 49, 50] and is a common technique in machine learning to improve the efficiency of computationally complex algorithms such as MCMC [50].
We briefly describe our approach to learning the time-shift distributions here but leave full technical details to Section 3 of S1 text. Let $\boldsymbol{\zeta}_{i}=(a_{i},b_{i},p_{i})$ denote the parameters of the time-shift distribution for some model parameters $\boldsymbol{\theta}_{i}$. We generate our dataset by sampling VL model parameters from a hyper-cube that encompasses our model priors. Our previous work [17] is used to determine the time-shift distribution parameters to give the paired data, $\{(\tilde{\boldsymbol{\theta}}_{i},\boldsymbol{\zeta}_{i})\}_{i=1}^{n}$. The time-shift distributions are checked for validity purposes since sampling on the hyper-cube does not necessarily produce feasible parameter combinations. Instead of naively sampling parameters, we check that the growth rate, $\lambda>0$, and that $95\%$ of the density for $\tau$ is between $[-7,7]$. The first condition ensures that there is a non-zero probability of the infection going non-extinct before reaching the exponential growth phase. The second condition ensures that long time-shifts, typically about a half of the total viral infection time-span (in our case), are considered un-biological and that time-shifts lying outside range likely suggests an incorrect model specification or poor prior choices. The goal here is to learn a function $g:\mathbb{R}^{5}\mapsto\mathbb{R}^{3}$ given by $g(\boldsymbol{\theta}_{i})=\boldsymbol{\zeta}_{i}$ to a high degree of accuracy. The learned function will approximate the mapping from the VL model parameters to the time-shift parameters without the need to perform complex numerical routines. Replacing the numerical routine with this learned function makes the runtime to find the parameters approximately 200$\times$ faster.
2.8.3 Software
@@ 2.9 Datasets
In this analysis, we use two datasets: a published dataset of COVID-19 VL measurements collected during the 2020-2021 NBA season restart, and a synthetic dataset designed to resemble the NBA dataset for method validation. The synthetic dataset serves as a controlled setting with frequent, low-noise observations and known parameters. It consists of $N=100$ individuals and is generated using our prior model for the hyperparameters $\boldsymbol{\phi}$. The full data generation process is detailed in Section 2.9.1, and the simulated cohort is shown in Fig. 3A.
@@
The limit of detection (LOD) in the NBA dataset is 2.658 log copies/mL and frequent testing and individuals experiencing only single infections resulted in many observations (i.e. some individuals having 80% LOD observations) at this threshold for each individual. Since it is most likely that individuals are only actually positive for some number of days pre/post the time where they have above LOD observations, we truncated the time series, retaining only the LOD measurements such that above-threshold observations remained between them. In some cases, multiple values below the LOD occur before another measurement slightly above the detection limit which complicates fitting. To address this, we further truncated time series at three consecutive threshold observations [6]. Finally, there were some short time-series—i.e. less than 2 observations above the LOD—and these were excluded. After filtering our dataset contains 163 of the original 241 timeseries. Note that we use the same post-processing procedure and LOD for the synthetic dataset. The NBA data is shown in Fig. 3B alongside our simulated cohort in panel Fig. 3A. The NBA data (Fig.3B) exhibits visibly higher variance across the entire time horizon compared to the simulated dataset (Fig.3A). This is primarily because the synthetic data were deliberately generated with lower noise to enable more controlled validation of our method.
@@
Our hierarchical model shares information from individuals with high information to improve estimates for those with sparse data, a process known as partial pooling [46, Chapter 5]. Previous analyses of this dataset [5, 12, 4, 6] used deterministic models [5, 12] or piecewise linear functions [12, 4, 6], whereas our model explicitly accounts for stochasticity in VL trajectories. Since we include individuals with both high and low information, direct parameter comparisons to Zitzmann et al. [5] are not possible. However, we can compare our $R_{0}$ estimates to their results and those from other SARS-CoV-2 studies [53] as well as other respiratory illnesses [31, 53]. Additionally, we compare predicted VL trajectories to those in Zitzmann et al. [5]. Phenomenological models [11, 4] estimate parameters based on piecewise linear models, which are not directly comparable to our mechanistic approach, but our model can estimate the same summary statistics as phenomenological models while also inferring biological quantities like $R_{0}$.
{adjustwidth}
-1in0in Refer to caption
@@
Fig 3: Scatter plots of the simulated and NBA datasets. Panel (a): the synthetic VL dataset generated according to the process in Section 2.9.1. Panel (b): the NBA VL dataset for the 2020-2021 seasons. In each panel an individual data point corresponds to a single measurement for a single individual. Points are shown as semi-transparent to approximate the “density” of points—highlighting outliers in panel (b). The red dashed line indicates the limit of detection.
2.9.1 Generating synthetic data
@@
We construct our synthetic dataset by initially sampling VL model parameters from the prior setup as outlined in Section 2.6. This requires fixing the shared parameters and sampling individual parameters from priors. The means of the shared parameters ($\mu_{R_{0}},\mu_{\delta},\mu_{\rho}$) are chosen to be in agreement with previous estimates [5]. The scale parameters ($\sigma_{R_{0}},\sigma_{\delta},\sigma_{\rho}$) were arbitrarily selected based on comparing the simulations to the NBA dataset. The measurement noise parameter $\kappa$ was chosen to be $0.5$ to reflect a reasonable amount of noise in the measured VL’s but still a “low noise” situation to assess the model in an ideal scenario. The values of all shared parameters are given in Table 3.
@@
Table 3: The values of the shared parameters used to generate the synthetic dataset.
Model parameters $\boldsymbol{\theta}_{i}$ are sampled for each individual from the priors assuming the values of the hyper-parameters in Table 3. Infection times are then drawn from the prior and then a time-series for each individual is simulated by running Algorithm 1. The variability in the priors causes a wide range of growth and decline rates as well as peak log-VL values which can cause unrealistic scenarios such as virus never being eliminated. In order to ensure we generate realistic data, we simulate under a rejection framework whereby we accept a simulation if an individual has peak log-VL between $5$ and $10$ (visually consistent with the range in the peak log-VL’s in the NBA dataset) and a corresponding peak infection time of between 4 and 9 days post-infection. We also require that the decline rate is consistent with a typical acute infection and so we require the log-VL to be less than $4.5$ at the end of the observation period (21 days post-infection). This condition ensures we do not have long-tails which are biologically invalid (i.e. no viral clearence). As in the NBA data, we assume a LOD of 2.658 log copies/mL, processing these as outlined in Section 2.9. Overall, the simulated dataset constitutes an idealised scenario where individuals have multiple observations in both the growth and decline phases.
@@ 3 Results
We apply our method to the two datasets outlined in Section 2.9. The initial exploration of the model fit to the synthetic dataset enables us to demonstrate the ability of the algorithm to produce accurate fits to the data for each individual in a situation where we know the true parameter values. The NBA dataset is used to demonstrate the utility of the method in a real-world scenario where the true parameter values are unknown.
All simulations and models were implemented in Julia [41] and run on a 2021 MacBook Pro with an M1 chip and 16 GB of RAM. Likelihood computations were parallelised across 4 high-performance CPU cores, as described in Section 2.8. A pilot run was used to tune the proposal distributions for both individual and shared parameters. Following this, we initialised four chains from perturbed versions of the parameters used to generate the synthetic data, to assess the robustness of the sampler to initial conditions. Each chain was run sequentially for 100,000 iterations, with the first 10,000 discarded as burn-in. The run times for the NBA dataset (the larger of the two datasets) took approximately 10 minutes per chain for a total runtime of approximately 45 minutes.
Trace plots for all shared parameters and a subset of individual parameters were visually inspected for divergences and showed no signs of poor mixing across the chains for both analyses. All parameters had $\hat{R}\leq 1.01$, indicating good within- and between-chain mixing [54], and effective sample sizes exceeded 1000 across both datasets. The acceptance rates were approximately $0.30$ and $0.4$ for the individual (averaged across groups) and shared parameters, respectively, both within the expected range for well-tuned random walk Metropolis proposals [55].
@@ 3.1 Simulation study
In this section, we present the inference results for the simulated dataset. The marginal distributions of the shared parameters, $\boldsymbol{\phi}$ and $\kappa$, are shown in Fig. 4, with prior distributions in yellow and posterior distributions in blue. True values, shown by the vertical dashed black lines, are computed from the sampled individual parameters since naively using the true parameters (i.e. Table 3) would be incorrect as our post-processing steps—like censoring observations—change how these are actually distributed. In what follows we refer to these estimated parameters as the “ground truth parameters”. For all parameters we observe substantial shifts from prior to posterior, indicating the influence of the data on the inference. All posteriors centered close to the ground truth values, demonstrating the method’s strength in estimating them.
{adjustwidth}
-2in0in Refer to caption
Fig 4: Density plots for the marginal distributions of the shared parameters (hyper-parameters and the scale of the measurement noise). Prior distributions are given by the orange lines and posteriors are represented by the blue histograms. The values indicated by the vertical dashed black lines are the true estimates of the parameters computed from the parameters used to simulate the dataset.
Fig. 5 shows the posterior predictive distributions for $R_{0}$, $\delta$, and $\rho$. These are obtained by using the posterior sampled hyper parameters and evaluating the normal model, i.e. for a draw $(\mu_{R_{0}},\sigma_{R_{0}})$, we sample $R_{0}^{\star}\sim\mathcal{N}(\mu_{R_{0}},\sigma_{R_{0}})$, and similarly for $\delta$ and $\rho$. The posterior predictive distributions represent the distribution of parameters that could be drawn if we sampled a new set of parameters. We can see from the posterior predictive distribution for all parameters that the model is able to appropriately capture the range of simulated parameters.
{adjustwidth}
-1.2in0in Refer to caption
Fig 5: Posterior Predictive Distributions for $R_{0}$, $\delta$, and $\rho$. The blue histograms show the posterior predictive distributions, while the grey shaded bars indicate the true parameter values used in the simulation. The grey histogram is noisy as there are only 100 parameter values sampled. The orange lines represent the prior predictive distributions. For clarity, we have zoomed in on the sampled values.
The difference between the posterior sampled parameters and the true values is shown in Fig. 6. Each subplot represents a different parameter, with the horizontal black line indicating a difference of 0 (strong parameter recovery). Intervals are coloured red if 0 is not within the 95% credible interval of the difference, and individual IDs are listed for ease of reference. In Fig. 6C, the true production rate, $\rho$, is contained within the 95% interval for all individuals except for individuals 1, 33 and 51. Fig. 6D shows that the infection times are generally well-estimated, with only 8 individuals not containing 0. This behaviour is expected due to the weakly informative priors on infection times and the censoring of data below the limit of detection that contributes to the slight misestimations. In the worst case, the infection times are off by up to 2 days. Fig. 6B shows 6 individuals where the interval for the difference in $\delta$ does not contain 0, suggesting these may be outliers. Similarly, in Fig 6A, 7 individuals have posteriors for $R_{0}$ that do not contain 0. These discrepancies are likely a result of pooling in our model, which will be further explored in the posterior predictive analysis shown next. Overall, even in cases where a parameter is slightly misestimated, the model still recovers the other parameters accurately.
{adjustwidth}
-2.0in0in Refer to caption
Fig 6: Interval plot showing the 95% intervals of the difference between posterior parameters and the true values for the simulated dataset. Panels (A)-(D) show the 95% interval of the difference between the posterior parameters and the true individual parameters. Parameter of interest is listed in the title of each subplot. The horizontal black dashed lines indicate a difference of 00. The white circles indicate the median difference. Individuals with 00 not in the 95% intervals have their intervals coloured in red and their ID is listed above or below the interval (depending on direction of difference).
Fig. 7 shows the posterior predictive fits for the synthetic dataset for the 22 individuals where a parameter was outside the 95% interval of the differences in Fig. 6. In all cases, the modelled trajectories track the observed data well. Cases where $\delta$ appears to be misestimated often coincide with visually noisier data in the decline phase (e.g., IDs 31, 44, 52, 59, 77, and 81). In these instances, most observations lie close to the median trajectory, with discrepancies likely arising due to noise in the data rather than model misspecification.
{adjustwidth}
-2.2in0in Refer to caption
@@
Fig 7: Posterior predictive fits for the synthetic dataset. Posterior predictive VL trajectories for individuals with 00 not in one of the intervals in Fig. 6. Solid lines represent the median of the posterior predictive fits and the shaded regions represent the 80% and 95% prediction intervals. Black dots indicate observed data. All subplots show the same time period $[-10,21]$.
@@
In addition to the trajectories and estimated parameters, by its construction, our model can estimate the posterior predictive time-shift for individuals. Here we show only a single individual as the other results all reflect the same prior predictive time-shift samples (orange) and similar posterior predictive curves (in blue). The time-shifts are shown in Fig. 8 where we see that the prior predictive time-shift distributions are quite wide and highly varied. Some of the density functions have large amounts of density concentrated above 1-2 days and have long tails (in both directions). Contrasting this, the posterior predictive time-shift distributions (blue) are more concentrated, with the distributions roughly resembling one another besides the variation arising from the posterior sampled parameters themselves.

### v2

@@ 2.7 Inference
The form of Eq. (5) suggests a Gibbs sampler that updates the individual-level parameters ($i=1,\dots,N$) and the shared parameters $(\boldsymbol{\phi},\kappa)$ via their conditionals. However, these conditionals are not available in closed form and cannot be sampled from directly, so we employ Metropolis-Hastings updates within a Gibbs sampling framework [48, Chapter 10]. Full algorithmic details of the sampler, including parallelisation strategies, are provided in Section 5 of S1 text..
@@ 2.8 Datasets
In this analysis, we use a collection of datasets: a published dataset of COVID-19 VL measurements collected during the 2020-2021 NBA season restart, and 50 synthetic datasets for method validation. Each synthetic dataset serves as a controlled setting with frequent, low-noise observations and known parameters. The purpose of the simulated datasets is to validate the method can recover parameters in an ideal setting. Each dataset consists of $N=100$ individual timeseries and are generated using our prior model for the hyperparameters $\boldsymbol{\phi}$. The full data generation process is detailed in Section 2.8.1, and an example simulated cohort is shown in Fig. 2A.
@@
The limit of detection (LOD) in the NBA dataset is 2.658 log copies/mL and frequent testing and individuals experiencing only single infections resulted in many observations (i.e. some individuals having 80% LOD observations) at this threshold for each individual. Since it is most likely that individuals are only actually positive for some number of days pre/post the time where they have above LOD observations, we truncated the time series, retaining only the LOD measurements such that above-threshold observations remained between them. In some cases, multiple values below the LOD occur before another measurement slightly above the detection limit which complicates fitting. To address this, we further truncated time series at three consecutive threshold observations. Finally, there were some short time-series—i.e. less than 2 observations above the LOD—and these were excluded. After filtering our dataset contains 163 of the original 241 timeseries. Note that we use the same post-processing procedure and LOD for the synthetic dataset. The NBA data is shown in Fig. 2B alongside an example simulated cohort in panel Fig. 2A. The NBA data (Fig. 2B) exhibits visibly higher variance across the entire time horizon compared to the simulated dataset (Fig. 2A). This is primarily because the synthetic data were deliberately generated with lower noise to enable more controlled validation of our method.
@@
Our hierarchical model shares information from individuals with high information to improve estimates for those with sparse data, a process known as partial pooling [51, Chapter 5]. Previous analyses of this dataset [5, 12, 4, 6] used deterministic models [5, 12] or piecewise linear functions [12, 4, 6], whereas our model explicitly accounts for stochasticity in VL trajectories. Since we include individuals with both high and low information, direct parameter comparisons to the model presented in the Supplementary Information (SI) of Zitzmann et al. [5] are not possible. However, we can compare our $R_{0}$ estimates to their results and those from other SARS-CoV-2 studies [52] as well as other respiratory illnesses [31, 52]. Additionally, we compare predicted VL trajectories to those of some of the SI fits in Zitzmann et al. [5]. Phenomenological models [11, 4] estimate parameters based on piecewise linear models, which are not directly comparable to our mechanistic approach, but our model can estimate the same summary statistics as phenomenological models while also inferring biological quantities like $R_{0}$.
@@
Fig 2: Scatter plots of an example simulated dataset and the NBA dataset. Panel (a): an example synthetic VL dataset generated according to the process in Section 2.8.1. Panel (b): the NBA VL dataset for the 2020-2021 seasons. In each panel an individual data point corresponds to a single measurement for a single individual. Points are shown as semi-transparent to approximate the “density” of points—highlighting outliers in panel (b). The red dashed line indicates the limit of detection.
2.8.1 Generating synthetic datasets
@@
We construct our synthetic datasets by initially sampling VL model parameters from the prior setup as outlined in Section 2.6. This requires fixing the shared parameters and sampling individual parameters from priors. The means of the shared parameters ($\mu_{R_{0}},\mu_{\delta},\mu_{\rho}$) are chosen to be in agreement with previous estimates [5, SI]. The scale parameters ($\sigma_{R_{0}},\sigma_{\delta},\sigma_{\rho}$) were arbitrarily selected based on comparing the simulations to the NBA dataset. The measurement noise parameter $\kappa$ was chosen to be $0.5$ to reflect a reasonable amount of noise in the measured VLs but still a “low noise” situation to assess the model in an ideal scenario. The values of all shared parameters are given in Table 3.
@@
Table 3: The values of the shared parameters used to generate the synthetic datasets.
Model parameters $\boldsymbol{\theta}_{i}$ are sampled for each individual from the priors assuming the values of the hyper-parameters in Table 3. Infection times are then drawn from the prior and then a time-series for each individual is simulated by running Algorithm 1. The variability in the priors causes a wide range of growth and decline rates as well as peak log-VL values which can cause unrealistic scenarios such as virus never being eliminated. In order to ensure we generate realistic data, we simulate under a rejection framework whereby we accept a simulation if an individual has peak log-VL between $5$ and $10$ (visually consistent with the range in the peak log-VLs in the NBA dataset) and a corresponding peak infection time of between 4 and 9 days post-infection. We also require that the decline rate is consistent with a typical acute infection and so we require the log-VL to be less than $4.5$ at the end of the observation period (21 days post-infection). This condition ensures we do not have long-tails which are biologically invalid (i.e. no viral clearence). As in the NBA data, we assume a LOD of 2.658 log copies/mL, processing these as outlined in Section 2.8. Overall, the simulated datasets constitute an idealised scenario where individuals have multiple observations in both the growth and decline phases.
@@ 3 Results
In this section we apply our method to the datasets outlined in Section 2.8. The initial exploration of the model fit to the synthetic datasets enables us to demonstrate the ability of the algorithm to produce accurate fits to the data in a situation where we know the true parameter values. The NBA dataset is then used to demonstrate the utility of the method in a real-world scenario where the true parameter values are unknown.
All simulations and models were implemented in Julia [42] and run on a 2021 MacBook Pro with an M1 chip and 16 GB of RAM. Likelihood computations were parallelised across 4 high-performance CPU cores, as described in Section 5 of S1 text.. Pilot runs were used to tune the proposal distributions for both individual and shared parameters. Following this, we initialised three chains from perturbed versions of the parameters used to generate the synthetic data, to assess the robustness of the sampler to initial conditions. Each of the chains were run for 100,000 iterations, with the first 20,000 discarded as burn-in. The run times for the NBA dataset (the larger of the datasets) took approximately 10 minutes per chain for a total runtime of approximately 30 minutes.
Trace plots for all shared parameters and a subset of individual parameters were visually inspected for divergences and showed no signs of poor mixing across the chains for both analyses. All parameters had $\hat{R}\leq 1.01$, indicating good within- and between-chain mixing [53], and minimum effective sample sizes exceeding 500. The acceptance rates were approximately $0.30$ for the individual (averaged across groups) and shared parameters, respectively, both within the expected range for well-tuned random walk Metropolis proposals [54].
@@ 3.1 Simulation study
Table 4 reports numerical summaries of the posterior fits, while Fig. 3 shows the 95% credible intervals for the population parameters. The true values fall within these intervals in most cases, with the lowest empirical coverage at 0.90. The credible intervals are narrow relative to the scale of the parameters (Interval width column in Table 4), indicating that the model recovers the population values with reasonable precision. Relative bias was computed as the bias divided by the true parameter value, giving a signed percentage difference. For all parameters except $\sigma_{R_{0}}$ and $\sigma_{\rho}$, posterior means are within $\pm 3.7\%$ of the true value. For $\sigma_{R_{0}}$ and $\sigma_{\rho}$ the relative biases are 0.114-0.114 and $0.626$, respectively. However, Fig. 3 shows that the 95% credible intervals for these variance parameters are much wider (average widths 1.036 and 0.827), reflecting substantial uncertainty. This explains why, despite noticeable bias in the posterior means of these two parameters, interval coverage remains adequate.
Parameter Coverage Interval width Relative bias
$\mu_{R_{0}}$ $0.90$ 0.964[0.901,1.051]0.964\>[0.901,1.051] 0.021[0.013,0.059]0.021\>[-0.013,0.059]
$\sigma_{R_{0}}$ $1.00$ 1.036[0.808,1.233]1.036\>[0.808,1.233] 0.114[0.397,0.23]-0.114\>[-0.397,0.23]
$\mu_{\delta}$ $0.920$ 0.086[0.081,0.092]0.086\>[0.081,0.092] 0.008[0.027,0.013]-0.008\>[-0.027,0.013]
$\sigma_{\delta}$ $0.90$ 0.056[0.052,0.06]0.056\>[0.052,0.06] 0.037[0.162,0.1]-0.037\>[-0.162,0.1]
$\mu_{\rho}$ $0.920$ 0.750[0.685,0.823]0.750\>[0.685,0.823] 0.017[0.053,0.083]0.017\>[-0.053,0.083]
$\sigma_{\rho}$ $0.980$ 0.827[0.665,1.015]0.827\>[0.665,1.015] 0.626[0.011,1.643]0.626\>[-0.011,1.643]
$\kappa$ $0.940$ 0.044[0.042,0.045]0.044\>[0.042,0.045] 0.006[0.025,0.027]0.006\>[-0.025,0.027]
Table 4: Numerical summary of posterior fits for the shared parameters. Coverage is the proportion of the 50 simulations in which the true parameter value lies within the 95% credible interval. The interval width is averaged across the 50 simulations, with an 80% confidence interval derived from quantiles ($q_{10},q_{90}$) reported. Relative bias is computed by dividing the bias by the true value to capture the scale of the difference, again with 80% confidence interval provided. An 80% confidence level was chosen to provide adequate precision given only 50 observations (datasets).
Fig 3: 95% credible interval plots of the population parameters for the simulated study. Blue lines indicate the 95% credible interval for a given simulation, with blue points indicating the median. The horizontal red dashed line shows the true value of the parameters used throughout the simulations.
Fig. 4 shows the posterior predictive fits for some example individuals from the synthetic dataset shown in Fig. 2. In all cases, the modelled trajectories track the observed data well, demonstrating that the method can reliably recover the viral dynamics in an ideal scenario.
@@
Fig 4: Posterior predictive fits for a subset of individuals in an example synthetic dataset. Posterior predictive VL trajectories for a subset of individuals shown in Fig. 2. Solid lines represent the median of the posterior predictive fits and the shaded regions represent the 80% and 95% prediction intervals. Black dots indicate observed data.
@@
In addition to the trajectories and estimated parameters, by its construction, our model can estimate the posterior predictive time-shift for individuals. Fig. 5 compares samples from the the prior (orange) and posterior (blue) time shift distributions for a single individual. We see that the prior predictive distributions are quite wide and highly varied. Some of the density functions have large amounts of mass concentrated above 1-2 days and have long tails (in both directions). Contrasting this, the posterior predictive time-shift distributions (blue) are more concentrated, with the distributions roughly resembling one another besides the variation arising from the posterior sampled parameters themselves.